# The long-term stability of the Vast Polar Structure and its connection to a possible previous passage of the LMC

### v1

@@ Abstract
Context. The Vast Polar Structure (VPOS) is a thin, planar arrangement of co-orbiting dwarf galaxies, nearly perpendicular to the Milky Way (MW) disc. Its origin and long-term evolution remain open questions.
@@
Methods. We identify VPOS member galaxies and integrate their orbits over the past 5 Gyr using accurate phase-space data and time-evolving gravitational potentials that account for the mutual interaction between the MW and the Large Magellanic Cloud (LMC). The reconstructed trajectories allow us to examine the long-term evolution of the VPOS.
Results. We identify 15 galaxies as members of the VPOS, including 9 MW and 6 LMC satellites. We find that the VPOS has remained a stable structure, maintaining a roughly constant thickness ($\sim$15 kpc), flattening ($c/a\sim 0.2$), and orientation over time. The VPOS is not a transient alignment, but a long-lived planar structure in the MW system, that has persisted for at least the last 5 Gyrs. Notably, the VPOS predates the LMC’s infall, ruling out formation scenarios tied to its recent approach to the MW. While the LMC exerts a strong gravitational influence on the MW satellites, its impact on the VPOS is limited, leading only to mild perturbations such as an elongation in the galaxy distribution or a temporary compression of the plane. The structural properties of the VPOS remain almost unchanged, whether or not LMC satellites are included in the analysis, indicating a smooth dynamical integration with the rest of VPOS members upon entering the MW virial radius. This minimal dynamical impact on the VPOS results from the remarkable alignment between the LMC’s orbit and the plane’s orientation.
Conclusions. The VPOS is a stable, long-lived structure that predates the recent infall of the LMC and retains nearly constant structural properties over the last 5 Gyrs. Our findings suggest a strong connection between the VPOS and the LMC, consistent with a scenario in which the LMC is on its second pericentre and the VPOS originated primarily from satellites stripped during the first passage.
@@ 1 Introduction
In the early stages of the study of the MW satellites, it was already noted that some of the dwarfs known at the time, in particular Draco, Sculptor and Ursa Minor, plus the Magellanic Stream and some globular clusters lied in a great circle structure perpendicular to the MW disc (Kunkel & Demers 1976; Lynden-Bell 1976). The presence of that alignment was repeatedly confirmed by the successive discoveries of new satellites following the same planar distribution (e.g. Kroupa et al. 2005; Metz et al. 2007). Known as the Vast Polar Structure (VPOS, Pawlowski et al. 2012b, 2013), such arrangement of satellites displays a thickness of 20-30 kpc, short-to-long axis ratio of $0.18-0.30$ and radius of $\sim 250$ kpc (Pawlowski 2018, 2021), with the majority of its members showing a coherent motion (Metz et al. 2008; Fritz et al. 2018; Li et al. 2021; Taibi et al. 2024). The VPOS is not a unique feature of the MW system. Other planes of satellites have also been reported around M31 (Metz et al. 2007; Ibata et al. 2013; Conn et al. 2013), and in galaxies beyond of the Local Group (LG), like M81 (Chiboucas et al. 2013), M101 (Merritt et al. 2014), Centaurus A (Tully et al. 2015) and NGC 253 (Martínez-Delgado et al. 2021), among others (see Pawlowski 2021).
Planes of satellites, and most notably the VPOS, have been a significant subject of debate during the last two decades. It has been argued that the likelihood of the existence of such arrangements is very low in the context of the currently most widely accepted cosmological framework, the $\Lambda$ Cold Dark Matter ($\Lambda$CDM) model. Kroupa et al. (2005) showed that the planar distribution of the 11 classical satellites of the MW is very unlikely should it had been drawn from an isotropic distribution. Many studies based on cosmological simulations have searched for similar anisotropic phase-space configurations, consistently finding that they are very rare (see Pawlowski 2018). A recent example is the case of the plane of satellites observed around NGC 4490, for which comparable structures were found in only 0.21% of analogous galaxies in the TNG50 cosmological simulation (Karachentsev & Kroupa, 2024; Pawlowski et al., 2024), highlighting how uncommon such configurations are in the $\Lambda$CDM model. This mismatch between predictions and observations is often referred to as the Planes of Satellite Galaxies Problem (Pawlowski, 2018). However, it is debated whether it is another possible $\Lambda$CDM small-scale problem (see Bullock & Boylan-Kolchin 2017) or if it does not challenge $\Lambda$CDM at all (e.g. Sawala et al. 2023).
@@
In order to better understand the origin of the VPOS and to assess the degree to which it challenges the $\Lambda$CDM model, it is essential to study its stability and persistence in time. Several works have tried to tackle this problem by integrating the orbits of the galaxies of the VPOS and then tracking the evolution of its thickness and axis ratios over time. Lipnicky & Chakrabarti (2017) made use of proper motions (PM) from the Hubble Space Telescope (HST) for the classical MW satellites and integrated their orbits in a static MW potential, finding that the significance of the plane vanished within a mere 1 Gyr. Similarly, Maji et al. (2017) performed forward integration of the orbits of the same 11 satellites to predict the future of the VPOS. They concluded that the VPOS would triple its width in 1 Gyr, and that ultimately the galaxies would abandon the plane. More recently, Sawala et al. (2023) has also argued that the plane is a transient structure, based on forward and backward integration using systemic PMs derived with Gaia Early Data Release 3 (EDR3, Gaia Collaboration et al. 2021). They found that the plane gets thicker both towards past and future times, and also that the orientation of the plane has changed $\sim 17$ degrees over the last 0.5 Gyrs.
@@ 2 Methods
To investigate the stability of the VPOS over time, we first need to identify which galaxies are currently part of it in order to backward integrate their orbits. The reconstructed orbits provide us with a record of the positions of these galaxies as a function of time and thus allow us to trace the evolution of the VPOS. In this Section we introduce the sample of galaxies to be studied, their phase-space coordinates, and the procedure used for the VPOS membership classification. We then outline the gravitational potentials used for the orbit integration, the integration procedure itself, and the plane fitting method used to quantify the structure’s stability and orientation over time. We note that we only present a brief overview of the galaxy sample, the potentials and the integration method, since full details will be available in Martínez-García et al. (in prep.).
@@ 2.1 Galaxy sample
Our study is based on a sample of 58 dwarf galaxies located within 500 kpc of the MW, for which 6D phase-space coordinates are available. The core of the sample is drawn from Battaglia et al. (2022), which provides accurate systemic PMs for a large number of LG dwarfs based on Gaia EDR3, alongside a compilation sky coordinates, distances, and radial velocities from the literature. We complement this core sample with additional systems, namely Aquarius III, Bootes V, Eridanus IV, Leo VI, Pegasus IV, the Sagittarius dSph, and the Small Magellanic Cloud (SMC), whose phase-space coordinates were obtained from the Local Volume Database (Pace, 2024). We note that the systemic PMs of the galaxies of the sample are primarily based on Gaia EDR3/DR3, except for the SMC, for which we adopt values from HST (Zivick et al., 2018). To ensure accurate orbit integration, we accounted for the effect of Gaia EDR3/DR3 systematics in the PMs using a procedure analogous to the one described in Battaglia et al. (2022), with the exception that we made use of the newly available QSO candidate table released with Gaia DR3 (qso_candidate, Gaia Collaboration et al. 2023).
@@ 2.2 VPOS membership selection
We determine which galaxies of the sample are members of the VPOS, following a procedure analogous to that of Taibi et al. (2024). First, we compute the specific angular momentum of each galaxy. To account for observational uncertainties, we perform a Monte Carlo (MC) scheme with $10^{3}$ iterations, that randomly samples the phase-space coordinates from normal distributions centered on their measured values, with standard deviations equal to their reported uncertainties. When uncertainties are asymmetric, we take as standard deviation their mean value. The sampled positions and velocities are then transformed into Galactocentric Cartesian coordinates with Astropy (Astropy Collaboration et al., 2013, 2018) using the following solar parameters: $R_{0}=8.122\pm 0.021$ kpc (GRAVITY Collaboration et al. 2018), $z_{\odot}=20.8\pm 0.3$ pc (Bennett & Bovy 2019), $V_{R,\odot}=-12.9\pm 3.0$ km s-1, $V_{\phi,\odot}=245.6\pm 1.4$ km s-1, and $V_{Z,\odot}=7.78\pm 0.09$ km s-1 (Drimmel & Poggio 2018). These parameters are also randomly sampled in each iteration. For each of these $10^{3}$ samples, we compute the orbital poles, i.e. the direction of the specific angular momentum vector, via the cross product of position and velocity, and transform them into Galactic coordinates. We then quantify the alignment of the orbital poles with the VPOS normal vector, which points towards $(l,b)=(169.3^{\circ},-2.8^{\circ})$ (Pawlowski & Kroupa, 2013; Fritz et al., 2018). For each galaxy, we compute the fraction of MC realizations whose orbital poles fall within the region defined by the area covering 10% of the sky around the VPOS normal vector direction (with an aperture of $36.87^{\circ}$), and its antipole. This fraction, $f_{\mathrm{VPOS}}$, is as a measure of the alignment with the VPOS. Galaxies with $f_{\mathrm{VPOS}}>0.95$ are classified as ’on-plane’, those with $f_{\mathrm{VPOS}}<0.05$ as ’off-plane’, and systems in the range $0.950.05<f_{\mathrm{VPOS}}<$ 0.95 are considered ’uncertain’.
@@ 2.3.2 Integration procedure
For orbit integration, we use the AGAMA software (Vasiliev, 2018, 2019) embedded within a MC scheme in order to account for the observational uncertainties. For each galaxy and potential, we run $10^{3}$ iterations, sampling phase-space coordinates from random normal distributions centered on their measured values and dispersion equal to their uncertainties. The sampled coordinates are then transformed into Galactocentric Cartesian coordinates. With these initial conditions, orbits are integrated backward using AGAMA’s orbit method in 1500 steps from $t=0$ to $t=-5$ Gyr (or $-4$ Gyr for the L2M10first potential). We limit the orbit integration to the past 5 Gyr, as MW-like halos are typically assembled by that epoch, with no major mergers occurring afterward (Santistevan et al., 2020; Sotillo-Ramos et al., 2022). Going further back in time would enter a regime where significant accretion events or massive mergers not captured by our time-evolving potentials are likely to affect the satellites’ dynamics, making the integration results increasingly unreliable. The output of the integration consists of a set of $10^{3}$ orbits per galaxy and potential.
@@ 3.1 Members of the VPOS
We studied the potential membership of 58 dwarf galaxies to the VPOS by analyzing the distribution of their orbital poles. For each galaxy, we computed the fraction of MC realizations, $f_{\mathrm{VPOS}}$, whose orbital poles lie within the 10% sky area centered on the VPOS normal direction and its antipode. This metric serves as a quantitative measure of alignment with the VPOS. The full list of studied galaxies, along with their membership classification, and $f_{\mathrm{VPOS}}$ are presented in Table 1. The distribution of the orbital poles in Galactic coordinates is shown in Figure 1. For each galaxy we represent the orbital poles from the MC realizations and their median values. Galaxies are marked according to their membership to the VPOS and they are split across panels based on their Galactocentric distance to ease visualization.
@@
Table 1: Studied galaxies and their membership to the VPOS. Column 1 shows the names of the galaxies, Column 2 their classification as on-plane (on), off-plane (off) or uncertain (?) members of the VPOS, and Column 3 the fraction of MC realizations where there orbital poles fall within the area encompassing the 10% of the sky around the VPOS normal direction and its antipole. Columns 4-6 follow the same structure.
@@
Figure 2: Spatial distribution and orbital histories of VPOS member galaxies. The panels show the positions of on-plane galaxies and their past trajectories under potential V23, at the present time (left) and at $t=-1$ Gyr (right). The coordinate system used ($x^{\prime},y^{\prime},z^{\prime}$) is not Galactocentric, but instead defined by the principal axes of the galaxy distribution at $t=0$; in this frame, the $z^{\prime}$-axis is aligned with the normal vector of the VPOS. Points indicate galaxy positions at the corresponding time, while lines trace their past trajectories. Red denotes MW satellites, blue indicates LMC satellites, and black marks the LMC itself. Carina II and Horologium II are shown in green. Although not formally classified as on-plane members, they are LMC satellites and thus likely part of the VPOS (see Section 3.1). Dashed lines are used to represent the trajectories of Bootes III and Sculptor, the two counter-rotating galaxies of the VPOS. A grey annulus at the origin represents the MW stellar disk for reference, with a radius of 15 kpc.
To further illustrate the spatial and kinematic coherence of the on-plane galaxies of the VPOS, in Figure 2 we represent their orbits. The orbits shown were integrated under the potential V23 using the nominal phase-space coordinates of each galaxy, without incorporating uncertainties. The figure shows face-on and edge-on views of the VPOS at the present time and at $t=-1$ Gyr, a moment chosen to better distinguish the MW and LMC satellite populations before they further approach and become difficult to disentangle. For context, Carina II and Horologium II are also shown but marked differently. The figure highlights the clear planar alignment and coherent motion of most on-plane satellites, with Bootes III and Sculptor standing out as notable counter-rotating outliers. Sculptor appears to come from large distances and has only recently completed a pericentre ($t\sim-0.5$ Gyr), a behavior also observed in potentials L3M10 and L3M11. In the other potentials, however, it completes multiple pericentres, resembling Bootes III, which has three in the past 5 Gyr. In parallel, the LMC and its associated satellite group clearly form a compact, dynamically coherent system following a trajectory closely aligned with the VPOS plane. In the earlier snapshot ($t=-1$ Gyr), the LMC group is seen approaching an already-established planar structure around the MW, suggesting that the plane of satellites predates the LMC’s recent infall and was not formed as a result of it.
@@
Finally, we explored how VPOS membership may have evolved over time by recalculating the orbital poles of each galaxy every 1 Gyr from their integrated orbits. As expected, the propagation of uncertainties in the integration cause the orbital poles to spread with time, making the $f_{\mathrm{VPOS}}>0.95$ criterion increasingly strict at earlier epochs. Despite this, some galaxies (e.g., Bootes III, Tucana IV or Ursa Minor) still meet it even at $t=-5$ Gyr. Most other on-plane galaxies retain their median orbital poles within or very near the VPOS area throughout the whole integration time. Some off-plane or uncertain members of the VPOS, like Draco II, Segue 1, and Tucana III, whose orbital poles lie in the border of the VPOS area or are partially inside at $t=0$, may also align with the VPOS at earlier times, what highlights that using strict boundaries to define the membership can obviate some possible members. A striking case is Grus II, which usually stays well outside the VPOS area for most of the integration time and only enters around $t\sim-0.5$ Gyr, becoming an on-plane galaxy. The timing of the realignment of its orbital pole, together with the analysis of its orbital history, suggests that it was likely caused by the recent close passage of the LMC, which reoriented Grus II’s trajectory (Martínez-García et al., in prep.). We thus conclude that the on-plane population of the VPOS has remained stable over the past 5 Gyr, with the notable exception of the recent inclusion of Grus II. To avoid potential bias in the study of the VPOS stability, we exclude Grus II from such analysis. Nonetheless, for clarity and consistency, we continue to use the term “VPOS members” or “on-plane galaxies” to refer to the remaining satellites.

### v2

@@ Abstract
Context. The Vast Polar Structure (VPOS) is a thin, planar arrangement of co-orbiting dwarf galaxies, nearly perpendicular to the Milky Way (MW) disc, whose origin and long-term evolution remain open questions.
@@
Methods. We identify VPOS member galaxies and integrate their orbits over the past 5 Gyr using accurate phase-space data and time-evolving gravitational potentials that account for the mutual interaction between the MW and the Large Magellanic Cloud (LMC).
Results. We identify 15 galaxies as members of the VPOS, including 9 MW and 6 LMC satellites. We find that the VPOS has remained a stable structure, maintaining a roughly constant thickness ($\sim$15 kpc), flattening ($c/a\sim 0.2$), and orientation over time. While the LMC exerts a strong gravitational influence on the MW satellites, its impact on the VPOS is limited, leading only to mild perturbations. The structural properties of the VPOS remain almost unchanged, whether or not LMC satellites are included in the analysis, indicating a smooth dynamical integration with the rest of VPOS members upon entering the MW virial radius. This minimal dynamical impact on the VPOS results from the remarkable alignment between the LMC’s orbit and the plane’s orientation.
Conclusions. The VPOS is not a transient alignment, but a long-lived planar structure in the MW system, that has persisted for at least the last 5 Gyrs. Notably, the VPOS predates the LMC’s infall, ruling out formation scenarios tied to its recent approach to the MW. Our findings suggest a strong connection between the VPOS and the LMC, consistent with a scenario in which the LMC is on its second pericentre and the VPOS originated primarily from satellites stripped during the first passage.
@@ 1 Introduction
In the early stages of the study of the MW satellites, it was already noted that some of the dwarfs known at the time, in particular Draco, Sculptor and Ursa Minor, plus the Magellanic Stream and some globular clusters lied in a great circle structure perpendicular to the MW disc (Kunkel & Demers 1976; Lynden-Bell 1976). The presence of that alignment was repeatedly confirmed by the successive discoveries of new satellites following the same planar distribution (e.g. Kroupa et al. 2005; Metz et al. 2007). Known as the Vast Polar Structure (VPOS, Pawlowski et al. 2012b, 2013), such arrangement of satellites displays a thickness of 20-30 kpc, short-to-long axis ratio of 0.180.300.18-0.30 and radius of $\sim 250$ kpc (Pawlowski 2018, 2021b), with the majority of its members showing a coherent motion (Metz et al. 2008; Fritz et al. 2018; Li et al. 2021; Taibi et al. 2024). The VPOS is not a unique feature of the MW system. Other planes of satellites have also been reported around M31 (Metz et al. 2007; Ibata et al. 2013; Conn et al. 2013), and in galaxies beyond of the Local Group (LG), like M81 (Chiboucas et al. 2013), M101 (Merritt et al. 2014), Centaurus A (Müller et al. 2018), NGC 253 (Mart´ınez-Delgado et al. 2021), and NGC 2750 (Paudel et al. 2021), among others (see Pawlowski 2021b).
[…]
@@
[…]
@@ 2 Methods
To investigate the stability of the VPOS over time, we first need to identify which galaxies are currently part of it in order to backward integrate their orbits. The reconstructed orbits provide us with a record of the positions of these galaxies as a function of time and thus allow us to trace the evolution of the VPOS. In this Section we introduce the sample of galaxies to be studied, their phase-space coordinates, and the procedure used for the VPOS membership classification. We then outline the gravitational potentials used for the orbit integration, the integration procedure itself, and the plane fitting method used to quantify the structure’s stability and orientation over time. We note that we only present a brief overview of the galaxy sample, the potentials and the integration method, since full details will be available in Martínez-García et al. (in prep.), a dedicated paper on the orbit integration of dwarf satellites around the MW.
@@ 2.1 Galaxy sample
Our study is based on a sample of 58 dwarf galaxies located within 500 kpc of the MW, for which 6D phase-space coordinates are available. The core of the sample is drawn from Battaglia et al. (2022), which provides accurate systemic PMs for a large number of LG dwarfs based on Gaia EDR3, alongside a compilation sky coordinates, distances, and radial velocities from the literature. We complement this core sample with additional systems, namely Aquarius III, Bootes V, Eridanus IV, Leo VI, Pegasus IV, the Sagittarius dSph, and the Small Magellanic Cloud (SMC), whose phase-space coordinates were obtained from the Local Volume Database (Pace, 2024). We note that the systemic PMs of the galaxies of the sample are primarily based on Gaia EDR3/DR3, except for the SMC, for which we adopt values from HST (Zivick et al., 2018). To ensure accurate orbit integration, we accounted for the effect of Gaia EDR3/DR3 systematics in the PMs using a procedure analogous to the one described in Battaglia et al. (2022), with the exception that we made use of the newly available QSO candidate table released with Gaia DR3 (qso_candidate, Gaia Collaboration 2023).
@@ 2.2 VPOS membership selection
We determine which galaxies of the sample are members of the VPOS, following a procedure analogous to that of Taibi et al. (2024). First, we compute the specific angular momentum of each galaxy. To account for observational uncertainties, we perform a Monte Carlo (MC) scheme with $10^{3}$ iterations, that randomly samples the phase-space coordinates from normal distributions centered on their measured values, with standard deviations equal to their reported uncertainties. When uncertainties are asymmetric, we take as standard deviation their mean value. The sampled positions and velocities are then transformed into Galactocentric Cartesian coordinates with Astropy (Astropy Collaboration 2013, 2018) using the following solar parameters: $R_{0}=8.122\pm 0.021$ kpc (distance from the Sun to the Galactic center; GRAVITY Collaboration 2018), $z_{\odot}=20.8\pm 0.3$ pc (height of the Sun above the Galactic midplane; Bennett & Bovy 2019), $V_{R,\odot}=-12.9\pm 3.0$ km s-1 (radial velocity of the Sun with respect to the Galactic center; positive values are directed outward from the center), $V_{\phi,\odot}=245.6\pm 1.4$ km s-1 (tangential velocity of the Sun in the direction of Galactic rotation), and $V_{Z,\odot}=7.78\pm 0.09$ km s-1 (vertical velocity of the Sun relative to the Galactic midplane, directed toward the North Galactic Pole; Drimmel & Poggio 2018). These parameters are also randomly sampled in each iteration. For each of these $10^{3}$ samples, we compute the orbital poles, i.e. the direction of the specific angular momentum vector, via the cross product of position and velocity, and transform them into Galactic coordinates. We then quantify the alignment of the orbital poles with the VPOS normal vector, which points towards $(l,b)=(169.3^{\circ},-2.8^{\circ})$ (Pawlowski & Kroupa, 2013; Fritz et al., 2018). For each galaxy, we compute the fraction of MC realizations whose orbital poles fall within the region defined by the area covering 10% of the sky around the VPOS normal vector direction (with an aperture of $36.87^{\circ}$), and its antipole. This fraction, $f_{\mathrm{VPOS}}$, is as a measure of the alignment with the VPOS. Galaxies with $f_{\mathrm{VPOS}}>0.95$ are classified as ’on-plane’, those with $f_{\mathrm{VPOS}}<0.05$ as ’off-plane’, and systems in the range $0.05<f_{\mathrm{VPOS}}<0.95$ are considered ’uncertain’.
@@ 2.3.2 Integration procedure
For orbit integration, we use the AGAMA software (Vasiliev, 2018, 2019) embedded within a MC scheme in order to account for the observational uncertainties. For each galaxy and potential, we run $10^{3}$ iterations, sampling phase-space coordinates from random normal distributions centered on their measured values and dispersion equal to their uncertainties. The sampled coordinates are then transformed into Galactocentric Cartesian coordinates. With these initial conditions, orbits are integrated backward using AGAMA’s orbit method in 1500 steps from $t=0$ to t=5t=-5 Gyr (or 4-4 Gyr for the L2M10first potential [^1] Potential L2M10first only allows us to integrate over the last 4 Gyrs, since the simulated time for such potential does not reach earlier times (see Vasiliev 2024).). We note that the effect of dynamical friction is not considered in the integration, however we do not expect it to have a significant impact in our analysis (we address the reader to Appendix A, where we further discuss this and other possible caveats of our work). We limit the orbit integration to the past 5 Gyr, as MW-like halos are typically assembled by that epoch, with no major mergers occurring afterward (Santistevan et al., 2020; Sotillo-Ramos et al., 2022). Going further back in time would enter a regime where significant accretion events or massive mergers not captured by our time-evolving potentials are likely to affect the satellites’ dynamics, making the integration results increasingly unreliable. The output of the integration consists of a set of $10^{3}$ orbits per galaxy and potential.
@@ 3.1 Members of the VPOS
We studied the potential membership of 58 dwarf galaxies to the VPOS by analyzing the distribution of their orbital poles. The full list of studied galaxies, along with their membership classification, and $f_{\mathrm{VPOS}}$ are presented in Table 2. The distribution of the orbital poles in Galactic coordinates is shown in Figure 1. For each galaxy we represent the orbital poles from the MC realizations and their median values. Galaxies are marked according to their membership to the VPOS and they are split across panels based on their Galactocentric distance to ease visualization.
Table 1: VPOS membership classification
@@
 [^2] Studied galaxies and their membership to the VPOS. Column 1 shows the names of the galaxies, Column 2 their classification as on-plane (on), off-plane (off) or uncertain (?) members of the VPOS, and Column 3 the fraction of MC realizations where there orbital poles fall within the area encompassing the 10% of the sky around the VPOS normal direction and its antipole. Columns 4-6 follow the same structure.
@@
Figure 2: Spatial distribution and orbital histories of VPOS member galaxies. The panels show the positions of on-plane galaxies and their past trajectories under potential V23, at the present time (left) and at t=1t=-1 Gyr (right). The coordinate system used ($x^{\prime},y^{\prime},z^{\prime}$) is not Galactocentric, but instead defined by the principal axes of the galaxy distribution at $t=0$; in this frame, the $z^{\prime}$-axis is aligned with the normal vector of the VPOS. Points indicate galaxy positions at the corresponding time, while lines trace their past trajectories. Red denotes MW satellites, blue indicates LMC satellites, and black marks the LMC itself. Carina II and Horologium II are shown in green. Although not formally classified as on-plane members, they are LMC satellites and thus likely part of the VPOS (see Section 3.1). Dashed lines are used to represent the trajectories of Bootes III and Sculptor, the two counter-rotating galaxies of the VPOS. A grey annulus at the origin represents the MW stellar disk for reference, with a radius of 15 kpc.
To further illustrate the spatial and kinematic coherence of the on-plane galaxies of the VPOS, in Figure 2 we represent their orbits. The orbits shown were integrated under the potential V23 using the nominal phase-space coordinates of each galaxy, without incorporating uncertainties. The figure shows face-on and edge-on views of the VPOS at the present time and at t=1t=-1 Gyr, a moment chosen to better distinguish the MW and LMC satellite populations before they further approach and become difficult to disentangle. For context, Carina II and Horologium II are also shown but marked differently. The figure highlights the clear planar alignment and coherent motion of most on-plane satellites, with Bootes III and Sculptor standing out as notable counter-rotating outliers. Sculptor appears to come from large distances and has only recently completed a pericentre ($t\sim-0.5$ Gyr), a behavior also observed in potentials L3M10 and L3M11. In the other potentials, however, it completes multiple pericentres, resembling Bootes III, which has three in the past 5 Gyr. In parallel, the LMC and its associated satellite group clearly form a compact, dynamically coherent system following a trajectory closely aligned with the VPOS plane. In the earlier snapshot (t=1t=-1 Gyr), the LMC group is seen approaching an already-established planar structure around the MW, suggesting that the plane of satellites predates the LMC’s recent infall and was not formed as a result of it.
@@
Finally, we explored how VPOS membership may have evolved over time by recalculating the orbital poles of each galaxy every 1 Gyr from their integrated orbits. As expected, the propagation of uncertainties in the integration cause the orbital poles to spread with time, making the $f_{\mathrm{VPOS}}>0.95$ criterion increasingly strict at earlier epochs. Despite this, some galaxies (e.g., Bootes III, Tucana IV or Ursa Minor) still meet it even at t=5t=-5 Gyr. Most other on-plane galaxies retain their median orbital poles within or very near the VPOS area throughout the whole integration time. Some off-plane or uncertain members of the VPOS, like Draco II, Segue 1, and Tucana III, whose orbital poles lie in the border of the VPOS area or are partially inside at $t=0$, may also align with the VPOS at earlier times, what highlights that using strict boundaries to define the membership can obviate some possible members. A striking case is Grus II, which usually stays well outside the VPOS area for most of the integration time and only enters around $t\sim-0.5$ Gyr, becoming an on-plane galaxy. The timing of the realignment of its orbital pole, together with the analysis of its orbital history, suggests that it was likely caused by the recent close passage of the LMC, which reoriented Grus II’s trajectory (Martínez-García et al., in prep.). We also investigated the converse scenario to the case of Grus II, namely, whether any galaxies that were originally consistent with the VPOS have been scattered out of it by the influence of the LMC, and found no evidence of such an event. We thus conclude that the on-plane population of the VPOS has remained stable over the past 5 Gyr, with the notable exception of the recent inclusion of Grus II. To avoid potential bias in the study of the VPOS stability, we exclude Grus II from such analysis. Nonetheless, for clarity and consistency, we continue to use the term “VPOS members” or “on-plane galaxies” to refer to the remaining satellites.